Describe how you would make 100ml of a 10mmol glucose standard
Answer #1

10 mmol of glucose = 10/1000 moles = .01 mole

molar mass of glucose = 180g/mol

mass of glucose required = .01 * 180 = 1.8 g

1.8 gram of solid glucose is weighed and then dissolve it in 100ml of water in a standard flask

this will give the required standard solution of glucose

concentration of this solution become = 0.1 M